OpenStudy (anonymous):

SOMEBODY HELP ME PLEASE ""What is wrong with the solution: x(x+3)=9 x=9 or x+3=9 x=6 So, the roots are 6 & 9"" i dont understand 

OpenStudy (anonymous):

Suppose \(x=9\) is a solution. Does the equation still hold? \[9(9+3)=9~~?\]

OpenStudy (anonymous):

The answer is no. You can only apply the step you used if the equation were equal to 0. What you should have done: \[x(x+3)-9=0\\ x^2+3x-9=0\] Then find the roots by factoring or applying the quadratic formula.

OpenStudy (acxbox22):

using the zero product rule the function has to be equal to zero in order to solve for example x(x+3)=0 x=0 and x=-3 x(x+3)=9<----not equal to zero

OpenStudy (anonymous):

Thank you to yall!!!!